T. Chemicals is a major producer of oil-based fertilizers in the US. The company’s stock is currently selling for $80 per share and there are 10 million shares outstanding. The company also has debt outstanding with a market value of $400 million. The interest rate on debt is 10%. The company’s current capital structure approximates well its target position. The company’s equity beta is equal to 2.0.

The company is considering an expansion project expected to generate a rate of return of 20% annually. Assuming a corporate tax rate of 50%, risk free rate of 8%, and the expected rate of return on the market portfolio of 17%, determine whether the company should go ahead with the project under the following two scenarios:

1. The project has the same risk level as the company.

2. The project’s risk is higher than that of the company. The project’s unlevered beta is 2.5. Also, consistent with its higher risk level, the project is expected to generate a rate of return of 25%. Further, because of the project’s higher risk level, the company has decided to use a more conservative capital structure represented by a debt-to-asset ratio of 15%.
Answer #1

Cost of equity for company = Risk free rate + Beta*(Market return-Risk free rate)

= 8%+2*(17%-8%) = 26%

Equity value = 10,000,000*80 = $ 800,000,000 = $ 800 million

Debt = $400 million

Total = 800+400 =$ 1200 million

Equity % = 800/1200 = 66.67%

Debt % = 33.33%

WACC = Equity %*Cost of equity + Debt %*Cost of Debt*(1-Tax rate)

=66.67%*26% + 33.33%*10%*(1-50%) = 19%

Since the Rate of return is greater than WACC, company should go for the project.

2. Unlevered Beta is used when project is based on 100% equity but above project has debt of 15% and equity of 85%.

Unlevered Beta = 2.5

Levered Beta = Unlevered Beta*(1+(1-Tax rate)*(debt/Equity))

= 2.5*(1+(1-50%)*(15%/85%)) = 2.721

Cost of equity = 8%+2.721*(17%-8%) =32.49%

WACC = 85%*32.49% + 15%*10%*(1-50%) = 28.37%

Required rate of return is 25%

Since Required rate of return is less than the cost of capital i.e. WACC, company should not go for the project